2. At timet 0, a 650.0 2 resistor is connected across the terminals of a 17.0 nF capacitor that carries a charge of 8.20 uc. a. What is the maximum current through the resistor? b. What is the current in the resistor at t 5.00 us? c. What is the charge remaining on the capacitor plates at t 5.00 us?
Answer #1


(a)

Imax = V/R = Q/CR = 8.2*10^-6/(17*10^-9*650) = 0.742 A       <<<<<---------answer


(b)


current decays as i = Imax*e^-(t/RC)

i = 0.742*e^-(5*10^-6/(650*17*10^-9))


i = 0.472 A   <<<<<---------answer

(c)


charge current decays as q = Q*e^-(t/RC)

i = 8.2*10^-6*e^-(5*10^-6/(650*17*10^-9))


i = 5.21 uC   <<<<<---------answer